QikTrix:

Complete the square to rewrite the following equation. Identify the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 4x + y2 − 6y = −4

crispyrat:

is there a screenshot? there is not enough info ;-;

loganisbored:

yes there is

crispyrat:

no not really for example x2 i don't know what that means, that could be a mistype of x*2 or x^2

crispyrat:

so im just gonna assum its x^2 x^2+4x+y^2-6y=4 the equation should be (a+b)(a+b)=a^2+2*a*b+b^2 so in this case we have to do that for both x^2+4x so half of 4 is 2 so the thing is (x+2)(x+2) if we get x^2+4x+4 so we add 4 to -4 then y^2 − 6y=(y-3)(y-3)= y^2 − 6y+9 so then we add the 9 to -4 as well so we have (x+2)^2+(y-3)^2=-4+4+9 (x+2)^2+(y-3)^2=9 Then the center is (-x,-y) and the radius is sqrt((-x)^2+(-y)^2) so using this what is the radius and the center ;-;

crispyrat:

@AZ can u check my work ;-;

@crispyrat wrote:
so im just gonna assum its x^2 x^2+4x+y^2-6y=4 the equation should be (a+b)(a+b)=a^2+2*a*b+b^2 so in this case we have to do that for both x^2+4x so half of 4 is 2 so the thing is (x+2)(x+2) if we get x^2+4x+4 so we add 4 to -4 then y^2 − 6y=(y-3)(y-3)= y^2 − 6y+9 so then we add the 9 to -4 as well so we have (x+2)^2+(y-3)^2=-4+4+9 (x+2)^2+(y-3)^2=9 Then the center is (-x,-y) and the radius is sqrt((-x)^2+(-y)^2) so using this what is the radius and the center ;-;

loganisbored:

your smart

crispyrat:

stop spamming. idrc what you say but don't spam on a question with unneeded compliments

loganisbored:

im not

loganisbored:

i said one thing and it was a compliment so chill out and stop complaining about one little thing

snowflake0531:

@crispyrat wrote:
so im just gonna assum its x^2 x^2+4x+y^2-6y=4 the equation should be (a+b)(a+b)=a^2+2*a*b+b^2 so in this case we have to do that for both x^2+4x so half of 4 is 2 so the thing is (x+2)(x+2) if we get x^2+4x+4 so we add 4 to -4 then y^2 − 6y=(y-3)(y-3)= y^2 − 6y+9 so then we add the 9 to -4 as well so we have (x+2)^2+(y-3)^2=-4+4+9 (x+2)^2+(y-3)^2=9 Then the center is (-x,-y) and the radius is sqrt((-x)^2+(-y)^2) so using this what is the radius and the center ;-;
You're all correct just one thing, radius is the r in \[(x-h)^2+(y-k)^2=r^2\] and the center is (h,k) I was a bit confused when you said "Then the center is (-x,-y) and the radius is sqrt((-x)^2+(-y)^2)"

snowflake0531:

Also, the standard form of the circle should be written in r^2 so \[(x+2)^2+(y-3)^2=3^2\]

crispyrat:

ah ok i must have misworded somethingsit was a long time since i used quadratics